# Metadata practices for simulation workflows

### v1

@@ 4.3 Hydrology use case
Here, we show how the proposed metadata managing guidelines and the Archivist can be applied to a workflow using a hydrologic model and which results can be derived from these (fig. 5). On the left, a hydrologic model is visually represented by the logo of the mesoscale Hydrologic Model (mHM [45, 46, 18]) in fig. 5A. Both, output variables and metadata information represented by the blue and yellow cylinder, respectively, are stored. In the hydrology use case, it is worth noting that the model output is orders of magnitude larger in terms of size than the associated metadata. The output variables are user-specified and contain a minimal set of metadata like, unit and creation date. The metadata managing guidelines presented here does also allow us to create a set of additional metadata that provide comprehensive information about the workflow at hand (yellow cylinder in fig. 5). For example, the entire configuration of the hydrologic model and its parameters can be recorded as well as the execution environment, version information, and all inputs and outputs. The Archivist is then able to process the metadata information in a user-specific way. Two potential applications of the Archivist are shown in fig. 5B and fig. 5C.
@@
In fig. 5B, a performance analysis of the hydrologic model is shown. A routine exercise in hydrologic modeling is parameter calibration. This stems from the fact that hydrologic modeling follows the paradigm that model parameters govern model behavior. For example, water infiltration into the soil is controlled by a shape parameter that dictates how fast soil infiltration is decreasing when the soil is drying. This flexibility is fundamental to account for highly conductive soils (i.e., sandy soils) and low conductive soils (i.e., clay soil). Here, we are able to analyze how model performance depends on the particular choice of a parameter (fig. 5B). This can be done in principle over all the simulations that have ever been done using the hydrologic model which is not possible without this comprehensive set of metadata. Performance is measured by comparing the model output, typically river discharge, with observations and calculate a performance measure from the simulated and observed time series. These performance measures are created in a way that the optimal value is at one and decreases to $\infty$- ∞ for simulated and observed time series that are unrelated [47]. Hydrologic models are run over a set of model parameters and given the performance of these, a cumulative distribution function can be calculated (fig. 5B)[18, 48]. fig. 5B shows that model configuration $\mathcal{P}_{1}$ is outperforming model configuration $\mathcal{P}_{2}$. Parameter set $\mathcal{P}_{1}$ leads to a larger frequency of performance around 0.5 and higher, while parameter set $\mathcal{P}_{2}$ has larger relative frequency for lower performance values of zero and less. It is important to highlight here that such an analysis can be done with all available parameter configurations, moreover subsets of parameters of interest can be created so that their performances are investigated. This allows users to obtain a deeper understanding of how combinations of model parameters are affecting model performance at a level of detail that is currently not possible.
@@
A critical task in the evaluation of hydrologic models is the visual inspection of the simulated and observed time series of a variable of interest. fig. 5C depicts two time series of river discharge. The time series are based on simulations (red line in fig. 5C) and the black line is based on observations. The visual inspection then allows understanding the impact of the parameter calibration. More precisely, it allows understanding which part of the hydrograph is simulated well (e.g., high river discharge, low river discharge values, or the transition phase between high and low values). Having an annotated database at hand, as shown in fig. 5, allows freely selecting parameter configurations from already created simulations, not necessarily related to parameter optimization. Ongoing discussions on modern hydrologic sciences raise the need for reusability of hydrologic models results to better understand how process parametrizations are affecting model performance and behavior [49]. Such efforts largely benefit from enriching model results with metadata describing the simulation experiment. For example discrepancies in model performance can be found to not only be due to differences in model parameters, but also model configuration.
@@
Figure 5: Hydrology use case. A) A hydrologic model, here represented by the logo of the mesoscale Hydrologic Model (mHM https://mhm-ufz.org, last access: 25 June 2024) creating output (i.e., hydrological variables shown by the blue cylinder) and additional metadata information (yellow cylinder). A hydrologic model simulates the water cycle at the land surface. A typical output variable is river discharge (m3s-1). Hydrological data is annotated by metadata post-processed by the Archivist. B) Distribution of prediction performances of the hydrologic model across measurement stations for two parameter sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Prediction performance is estimated by comparing simulated and observed river discharge for each measurement station. C) Time series of observed (black) and predicted river discharge (red; parameter set $\mathcal{P}_{1}$).
5 Discussion
@@
Motivated by the expectation that rigorous metadata management is the key to enhance reproducibility and interpretability of scientific results, this work proposes and applies practices for handling metadata in simulation research. We derived a generic knowledge production workflow to illustrate that different types of metadata can be collected at different steps of a simulation, then be post-processed, and finally exploited alongside data To facilitate post-processing, we developed he Archivist framework, a Python-based tool for parsing raw metadata files and combining the extracted information into a structured file. As example implementations, we presented a conceptual example and two use cases with varying degrees of complexity. First was a hypothetical minimal example consisting of simple generic time-driven simulations where the collected metadata was used to track the performance of the simulation measured by the real time factor, and the accuracy of the model obtained by analyzing the generated results. Even if simple in nature, processing the metadata stemming from these simulations can prove challenging when dealing with larger volumes. As the goal of the experiment was to find a compromise between performance and accuracy by exploring the parameter space, a consequently large amount of data and metadata was generated. By automating metadata processing and structuring with the Archivist, the experiment pipeline was streamlined from data generation to data usage. Our second use case was a proof of concept of a benchmarking and verification workflow. By running several simulations of the same model[38] with NEST GPU[36], the simulation performance was compared and simulation results were statistically verified across multiple hardware platforms. Although the software setup was similar, the format of the information retrieved directly from the hardware varied between platforms. By leveraging the modularity of the Archivist, parsing functions specific to each platform were implemented and could be exchanged without needing to modify the experiment workflow. As the parsing output was unified through a common structure, the processed metadata was homogenized across hardware platforms and was transparently exploited. Our last use case was a routine parameter calibration procedure where multiple parameters for the mesoscale Hydrologic Model (mHM [45, 46, 18]) are evaluated across several simulation configurations. Due to its routine nature, data and metadata for each procedure have accumulated over time. As information from each parameter set is collected, the collection of raw metadata is large even within a single calibration procedure. By filtering the metadata entries to specific parameter sets, the Archivist was used to efficiently explore the available database of results.
@@
With the Archivist’s modular design shown in “Metadata post-processing framework – the Archivist”, users only need to define parsing functions to extract information from metadata entries, and use a schema to combine the results into desirable structures and formats. Both parsing functions and schemas can be shared among groups to foster reusability of implemented workflows. However, we note the need to properly document the defined functions and structures. The same information can be interpreted differently by multiple users; mistakenly one might extract information with a different interpretation than expected. A simple example would be reading a number with a large amount of decimal digits. Functions can be implemented to use the exact number with all decimals, truncate after an arbitrary significance of decimal digits, or round to the nearest integer. All of these implementations are valid interpretations and their use depends on how the information will be exploited later. Appropriate documentation solves this problem by disambiguating between implementations and increasing interpretability of the processed metadata. Much like the source code and version of used models and simulators help describing a simulation experiment, the implementation and version of parsing functions describe the post-processing operations. This is of particular importance considering that metadata processing for a given dataset is not a single final process, but can be refined for different needs. By keeping a copy and sharing the raw metadata set, data can be subsequently annotated by the same or other users even if certain metadata entries were filtered out after an initial processing and annotation step. Data is then further enriched by extending the available annotated information and can later be used for new studies. Future contributions to the Archivist could exploit the formalization of datasets structures provided by the data model to automatically define parsing targets without the need of user definition. Furthermore, given a detailed enough description of dataset contents, parsing function could be automatically generated too.
@@
As mentioned in “Metadata post-processing framework – the Archivist”, there exist alternative tools for metadata post-processing, such as the AiiDA[15] parsers or DataLad[17] metadata extractors. AiiDA is a workflow manager that can use parsing plugins to extract information from files generated during the simulation workflow. The parsed information can then be used as an output of a workflow step or attached as additional metadata for data provenance tracking. DataLad is a distributed data management system where users can define extractors to parse files present in the database and extract metadata to annotate respective datasets based on the data versioning system git-annex (https://git-annex.branchable.com, last access: 01 March 2024). Conceptually, both these tools perform similar operations to the Archivist: users define functions to extract information from files, these functions are used to parse the contents of the generated or stored data and extract metadata according to specific needs, and for further reusability, users can share these functions. The difference lies in how parsing is orchestrated and how the extracted metadata is handled. In both AiiDA and DataLad, parsing is triggered on a file-by-file basis. Although the metadata can be integrated in the knowledge base (i.e., the provenance graph for AiiDA or as annotated information for DataLad), each piece of parsed information is treated independently and is not meant to be combined into a comprehensive output. In contrast, the parsers employed by the Archivist operate on an arbitrary number of files automatically determined by user defined rules. Furthermore, the Archivist permits the use of a custom JSON Schema to structure and combine the parsed metadata, thereby fostering interoperability and reusability of the post-processing results. We note that our intention is not to create a replacement for features in AiiDA or DataLad, but to implement a standalone tool that can transparently suit specific user needs. Conversely, dedicated workflow manager tools such as Snakemake[16] or AiiDA are particularly useful for implementing a complete new workflow. In addition to increased reusability and interoperability of workflows, these tools offer software environment handling, HPC cluster job handling, and data provenance tracking among other features[50]. Data provenance in particular is useful for describing the contents of the data generated during simulations. Indeed, these solutions are not incompatible with our framework but could actually synergize by using the Archivist as a parsing and structuring backend. To this end, future work can be done on a combined implementation of the Archivist with existing tools for automated provenance tracking and enhanced storage platform. This combined implementation would feature automated experiment description, flexible data characterization, and efficient data annotation and exploration methods. Finally, further parsing and structuring methods can be added to the Archivist by the user community driven by their specific needs. Sharing these methods in a common repository would pave the way for interoperable metadata descriptions.
6 Conclusion
@@
Our proposed practices for handling metadata in simulation workflows are applicable to a wide range of scientific disciplines with domain-specific workflows and metadata conventions. New workflows can benefit from these practices to ensure that metadata is handled accordingly. The practices can also be implemented in existing workflows without major restructuring of established code. In this case, our proof of concept tool, the Archivist, provides a flexible solution for parsing and structuring heterogeneous metadata files. Our practices and tool support sustainable numerical workflows, facilitating reproducibility and data reuse in generic simulation-based research.
7 Methods
7.1 Details on the neuroscience use case
@@
The neuroscience use case focuses on a workflow for benchmarking and verification of a spiking neural network simulator across multiple hardware platforms. The workflow implementation can be found at https://doi.org/10.5281/zenodo.13585723 (last access: 30 August 2024).
NEST GPU
@@
Hardware platforms
The interest on this experiment lies on asserting whether the simulator would produce the same results across different hardware platforms. As the models simulated rely on floating point calculations, artifacts during numerical computations may arise during simulation and could potentially accumulate to observable differences in produced results. In particular, GPU architectures can vary highly between generations, which increases the chances of such divergence. For this reason we test the simulator with different GPUs models, both consumer and data center level, with varying architectures: one laptop with the consumer GPU RTX 3070 Ti Mobile with CUDA version 12.2; two compute clusters, JUSUF [51] and JURECA-DC [52], both using CUDA version 11.3 and equipped with the data center GPUs V100 and A100, respectively; and a workstation with the consumer GPU RTX 4090 with CUDA version 11.4.
@@
7.2 Details on the hydroscience use case
@@
The hydroscience use case (“Hydrology use case”) presents how the metadata Archivist can be applied in the hydrological modeling sciences.
mHM
@@
An essential tool of this use case is the employed mesoscale hydrologic model mHM [45, 46, 18] (https://mhm-ufz.org, last access: 25 June 2024). mHM is a grid-based, spatially distributed hydrological model driven by daily precipitation, temperature, and potential evapotranspiration. It accounts for major hydrological processes such as snow accumulation and melt, canopy interception, soil infiltration, evapotranspiration, deep percolation, baseflow generation, and river routing. The open-source model code repository is available and is under active development and maintenance (https://git.ufz.de/mhm/mhm, last access: 12 February 2024). A general overview on the model processes and parameterization can be obtained from [45] and [46]. The model is an integral part of the German Drought Monitor (https://www.ufz.de/duerremonitor, last access: 25 June 2024). mHM was also applied and evaluated in multiple climatological regions, including Europe [55, 56], West Africa[57], India[58], and the conterminous United States[59, 48].
Hardware platforms
@@
mHM simulations were carried out on the High-Performance Computing (HPC) Cluster EVE, a joint effort of both the Helmholtz Centre for Environmental Research - UFZ (http://www.ufz.de/, last access: 25 June 2024) and the German Centre for Integrative Biodiversity Research (iDiv) Halle-Jena-Leipzig (https://www.idiv.de/, last access: 25 June 2024). The main compute hardware of EVE comprises a) 42 compute nodes with dual socket Intel Xeon 6348 CPUs with 512 Gigabytes of DDR4 main memory, two of these also include NVIDIA Tesla A100 GPGPUs, and b) 27 compute nodes with dual socket Intel Xeon Gold 6148 CPUs with up to 1536 Gigabytes of DDR4 main memory, two of which include NVIDIA Tesla V100 GPGPUs. The central network component of the cluster is an Intel Omni-Path 100 Series high performance interconnect, providing all compute nodes with non-blocking EDR bandwidth (100 Gigabit per second). All compute nodes share a 4.5 Petabyte IBM Spectrum Scale file system. The system performed with over 164 teraFLOPS under a High-Performance LINPACK (HPL) benchmark. For the simulations of this project, we have used CPU cores exclusively.
Model setup
@@
The mHM model is executed within the test basin provided along with the model source code, the so called "test basin". The test basin coming with the mHM source code is for the Moselle River basin upstream of Perl, a place, where the Moselle River leaves France and enters Luxembourg and Germany (Moselle Basin). The catchment area is approximately 11,500 $11,500\text{ km}^{2}$, altitude ranges between 150 and 1300 m. a.m.s.l. The Moselle River originates from the Vosges Mountains and is a tributary of the Rhine River. The origin of data used in the test example is provided on https://mhm-ufz.org/docs/ (last access: July 18th, 2024).
Performance metrics
@@
The hydroscience use cases makes use of performance metrics that are commonly used in hydrologic modeling. Explicitly, it is the Kling-Gupta efficiency (KGE [47]), that can be used to compare two time series. The metric combines the long-term mean, long-term variance, and Pearson correlation coefficient.
Data generation
@@
8 Acknowledgments

### v2

@@ Results
Here, we show how the proposed metadata management practices and the Archivist can be applied to a workflow using a hydrologic model and which results can be derived from these (fig. 5). A hydrologic model is visually represented by the logo of the mesoscale Hydrologic Model (mHM [47, 48, 20]) on the left in fig. 5A. Both output variables and metadata information, represented by the blue and yellow cylinder, respectively, are stored. In the hydrology use case, it is worth noting that the model output is orders of magnitude larger in terms of size than the associated metadata. The output variables are user-specified and contain a minimal set of metadata like unit and creation date. The practices presented here also allow us to create a set of additional metadata that provide comprehensive information about the workflow at hand (yellow cylinder in fig. 5). For example, the entire configuration of the hydrologic model and its parameters can be recorded as well as the execution environment, version information, and all inputs and outputs. The Archivist is then able to process the metadata information in a user-specific way. Two potential applications of the Archivist are shown in fig. 5B and fig. 5C.
@@
In fig. 5B, a performance analysis of the hydrologic model is shown. A routine exercise in hydrologic modeling is parameter calibration. This stems from the fact that hydrologic modeling follows the paradigm that model parameters govern model behavior. For example, water infiltration into the soil is controlled by a shape parameter that dictates how fast soil infiltration is decreasing when the soil is drying. This flexibility is fundamental to account for highly conductive soils (i.e., sandy soils) and low conductive soils (i.e., clay soil). Here, we are able to analyze how model performance depends on the particular choice of a parameter (fig. 5B). This can be done in principle over all the simulations that have ever been done using the hydrologic model, which is not possible without this comprehensive set of metadata. Performance is measured by comparing the model output, typically river discharge, with observations and calculating a performance measure from the simulated and observed time series. These performance measures are created in a way that the optimal value is at one and decreases to $\infty$- ∞ for simulated and observed time series that are unrelated [49]. Hydrologic models are run over a set of model parameters and given the performance of these, a cumulative distribution function can be calculated (fig. 5B)[20, 50]. fig. 5B shows that model configuration $\mathcal{P}_{1}$ is outperforming model configuration $\mathcal{P}_{2}$. Parameter set $\mathcal{P}_{1}$ leads to a larger frequency of performance around 0.5 and higher, while parameter set $\mathcal{P}_{2}$ has larger relative frequency for lower performance values of zero and less. Such an analysis can be done with all available parameter configurations, moreover subsets of parameters of interest can be created so that their performances are investigated. This allows users to obtain a deeper understanding of how combinations of model parameters are affecting model performance at a level of detail that is currently not possible.
@@
A critical task in the evaluation of hydrologic models is the visual inspection of the simulated and observed time series of a variable of interest. fig. 5C depicts two time series of river discharge. The time series are based on simulations (red line in fig. 5C) and the black line is based on observations. The visual inspection then allows an understanding of the impact of the parameter calibration. More precisely, it allows an understanding of which part of the hydrograph is simulated well (e.g., high river discharge values, low river discharge values, or the transition phase between high and low values). Having an annotated database at hand, as shown in fig. 5, allows users to freely select parameter configurations from already created simulations, not necessarily related to parameter optimization. Ongoing discussions on modern hydrologic sciences raise the need for reusability of hydrologic models results to better understand how process parametrizations are affecting model performance and behavior [51]. Such efforts largely benefit from enriching model results with metadata describing the simulation experiment. For example, discrepancies in model performance can be found to not only be due to differences in model parameters, but also model configuration.
@@
Figure 5: Hydrology use case. A) A hydrologic model, here represented by the logo of the mesoscale Hydrologic Model (mHM https://mhm-ufz.org, last access: 25 June 2024), creates output (i.e., hydrological variables shown by the blue cylinder) and additional metadata information (yellow cylinder). A hydrologic model simulates the water cycle at the land surface. A typical output variable is river discharge (m3s-1). Hydrological data is annotated by metadata post-processed by the Archivist. B) Distribution of prediction performances of the hydrologic model across measurement stations for two parameter sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Prediction performance is estimated by comparing simulated and observed river discharge for each measurement station. C) Time series of observed (black) and predicted river discharge (red; parameter set $\mathcal{P}_{1}$).
Discussion
@@
Motivated by the expectation that rigorous metadata management is a prerequisite of reproducibility and interpretability of scientific results, this work proposes and applies practices for handling metadata in simulation research. We derived a generic knowledge production workflow to illustrate that different types of metadata can be collected at different steps of a simulation, then be post-processed, and finally exploited alongside data. To facilitate post-processing, we developed the Archivist framework, a Python-based tool for parsing raw metadata files and combining the extracted information into a structured file. The primary purpose of the tool is to attach it to existing workflows that require routine execution and do not incorporate other means of metadata tracking functionality yet. As example implementations, we presented a conceptual example and two use cases with varying degrees of complexity. Our first use case was a hypothetical minimal example consisting of simple time-driven simulations where the collected metadata was used to track the performance of the simulation measured by the real time factor and the accuracy of the model obtained by analyzing the generated results. Even if processing a single metadata entry is simple in nature, processing large volumes of metadata stemming from these simulations can prove challenging . As the goal of the experiment was to find a compromise between performance and accuracy by exploring the parameter space, a consequently large amount of data and metadata was generated. By automating metadata processing and structuring with the Archivist, the experiment pipeline, from data generation to data usage, was streamlined. Our second use case was a proof of concept of a benchmarking and verification workflow. By running several simulations of the same model[40] with NEST GPU[38], the simulation performance was compared and simulation results were statistically verified across multiple hardware platforms. Although the software setup was similar on all platforms, the format of the information retrieved directly from the hardware varied . By leveraging the modularity of the Archivist, parsing functions specific to each platform were implemented and could be exchanged without needing to modify the experiment workflow. Because we uniformly structured the parsing output, the processed metadata describing different hardware platforms was homogenized and could be accessed the same way. Our last use case was a routine parameter calibration procedure where multiple parameters for the mesoscale Hydrologic Model (mHM [47, 48, 20]) were evaluated across several simulation configurations. Due to its routine nature, data and metadata for each procedure had accumulated over time. Furthermore, the collection of raw metadata is large even within a single calibration procedure, because information from each parameter set is collected. The Archivist helped combine and structure the metadata entries, allowing them to be stored in a database which could be queried for specific parameter sets to efficiently explore the results.
@@
The current version of the Archivist helps access and structure heterogeneous metadata collected from simulation workflows, yet it has limitations and room for improvement. With the Archivist’s modular design, described in “Metadata post-processing framework – the Archivist”, users only need to define parsing functions to extract information from metadata entries, and use a data template to combine the results into desirable structures and formats. Both parsing functions and data templates can be shared among groups to foster reusability of implemented workflows. However, we note the need to properly document the defined functions and structures. The same information can be interpreted differently by multiple users; one might mistakenly extract information with a different interpretation than expected. A simple example would be reading a number with a large amount of decimal digits. Functions can be implemented to use the exact number with all decimals, truncate after an arbitrary significance of decimal digits, or round to the nearest integer. All of these implementations are valid interpretations and their use depends on how the information will be exploited later. Appropriate documentation solves this problem by disambiguating implementations and increasing interpretability of the processed metadata. Much like the source code and version of used models and simulators help describe a simulation experiment, the implementation and version of parsing functions describe the post-processing operations. This is of particular importance considering that metadata processing for a given dataset is not a single final process, but can be refined for different needs. By keeping a copy and sharing the raw metadata set, data can be subsequently annotated by the same or other users even if certain metadata entries were filtered out after an initial processing and annotation step. Data is then further enriched by extending the available annotated information and can later be used for new studies. Future contributions to the Archivist could automatically interpret information in a data model to deduce the locations of files in a dataset and generate file target rules without the need for user definition Given a sufficiently detailed description of dataset contents, parsing functions could also potentially be generated.
@@
As mentioned in the Section titled “Metadata post-processing framework – the Archivist”, there exist alternative tools for metadata post-processing, such as the AiiDA[15] parsers or DataLad[17] metadata extractors. AiiDA is a workflow manager that can use parsing plugins to extract information from files generated during the simulation workflow. The parsed information can then be used as an output of a workflow step or attached as additional metadata for data provenance tracking. DataLad is a distributed data management system where users can define extractors to parse files present in the database and extract metadata to annotate their respective datasets based on the data versioning system git-annex (https://git-annex.branchable.com, last access: 01 March 2024). Conceptually, both these tools perform similar operations to the Archivist: users define functions to extract information from files, these functions are used to parse the contents of the generated or stored data and extract metadata according to specific needs, and users can share these functions for further reusability. The difference lies in how parsing is orchestrated and how the extracted metadata is handled. In both AiiDA and DataLad, parsing is triggered on a file-by-file basis. Although the metadata can be integrated into a knowledge base (i.e., the provenance graph for AiiDA or as annotated information for DataLad), each piece of parsed information is treated independently and further processing is required to combine it into a comprehensive output. In contrast, the parsers employed by the Archivist operate on an arbitrary number of files automatically determined by user-defined rules. Furthermore, the Archivist empowers users with the possibility to define a custom data template based on JSON Schema to structure and combine the parsed metadata. The ability to use a custom data template helps ensure that previous and future post-processing results are consistently structured. Because our practices promote storing unprocessed workflow results as well as processed results, previous results can be reprocessed with a newer data template if a different structure is needed for a new purpose, thereby fostering reusability. Additionally, users can base their template on vocabularies or schemas such as RO-Crate[18], CodeMeta, or Bioschemas[19] for further interoperability of their post-processing results. We note that our intention is not to create a replacement for features in AiiDA or DataLad, but to implement a standalone tool that can transparently suit specific user needs. Dedicated workflow manager tools such as Snakemake[16] or AiiDA are particularly useful for implementing a completely new workflow - in addition to increased reusability and interoperability of workflows, these tools offer software environment handling, HPC cluster job handling, and data provenance tracking among other features[52]. Data provenance in particular is useful for describing the contents of data generated during simulations. Rather than being incompatible with our framework, these alternative tools could actually be combined with it by using the Archivist as a parsing and structuring backend. This combined implementation would feature automated experiment description, flexible data characterization, and efficient data annotation and exploration. To this end, future work could be done on a combined implementation of the Archivist with existing tools for automated provenance tracking and enhanced storage platforms.
We note that, although our abstract simulation workflow fig. 1 (blue) can represent a wide range of real-world workflows, it is certainly not an absolute representation. This implies that our metadata management practices defined through this abstract workflow are not fully applicable to all workflows. Very intricate workflows, for example, may deal with challenges that do not allow for a one-off intervention of incorporating the Archivist, but call for further customized solutions instead. Particular examples are highly evolving workflows containing steps that are frequently swapped, added, or removed requires keeping track of which metadata to collect and updating the corresponding post-processing methods. Even updating only the simulation workflow is time-consuming, so updating the metadata post-processing workflow too will require even more effort. Furthermore, simulation workflows that require complex dynamical inputs prepared in a separate preceding workflow, or that incorporate separate subsequent post-processing workflows – if metadata on these inputs or subsequent workflows is not collected during execution, they would be irretrievably lost. A final example where our practices may not be fully applicable are simulations where the size of produced data and/or metadata is too large to store both in “raw” and processed format. These workflows require that data and metadata are processed, possibly filtered, and annotated during runtime before being able to be stored. Decoupling the post-processing of metadata describing data generated during the simulation experiment is not beneficial in these cases and would most likely lead to inconsistency issues. Nonetheless, even if not fully applicable, our practices can serve as a stepping stone for these complex workflows as metadata describing scientific experiments remain essential for long-term storage and sharing results.
@@
Our proposed practices for handling metadata in simulation workflows are applicable to a wide range of scientific disciplines with domain-specific workflows and metadata conventions. New workflows can benefit from these practices to ensure that metadata is handled accordingly. The practices can also be implemented in existing workflows without major restructuring of established code. In this case, our proof of concept tool, the Archivist, provides a flexible solution for parsing and structuring heterogeneous metadata files. Through the use cases presented in this work we illustrate how our practices and tool support sustainable numerical workflows, fostering reproducibility and data reuse in simulation-based research.
Methods
Details on the neuroscience use case
@@ Methods
The “Neuroscience use case” focuses on a workflow for benchmarking and verification of a spiking neural network simulator across multiple hardware platforms. The workflow implementation can be found at https://doi.org/10.5281/zenodo.13585723 (last access: 30 August 2024).
NEST GPU.
@@
Hardware platforms.
The goal of this experiment is to determine whether the simulator would produce the same results across different hardware platforms. As the models simulated rely on floating point calculations, artifacts during numerical computations may arise during simulation and could potentially accumulate to observable differences in produced results. In particular, GPU architectures can vary highly between generations, which increases the chances of such divergence. For this reason we test the simulator with different GPUs models, both consumer and data center level, with varying architectures: one laptop with the consumer GPU RTX 3070 Ti Mobile with CUDA version 12.2; two compute clusters, JUSUF [53] and JURECA-DC [54], both using CUDA version 11.3 and equipped with the data center GPUs V100 and A100, respectively; and a workstation with the consumer GPU RTX 4090 with CUDA version 11.4.
@@
Vocabulary usage example.
In Listing LABEL:lst:vocab_example we show a partial example output of a metadata post-processing pipeline. In this example, input and output file names were extracted from the workflow configuration file and formatted according to the Bioschemas vocabulary (https://bioschemas.org/types/ComputationalWorkflow/1.0-RELEASE, last visited: 23 January 2025). If specific file names are changed in the configuration this update would be reflected in the output of a subsequent metadata post-processing pipeline. These were then used to complete a simple Bioschemas document bioschemas.jsonld included with the workflow source code. Although valid, the document content is minimal and only shown as a proof of feasibility. Users would naturally employ the full capabilities of Bioschemas to describe their own workflows.
"inputs": [
"@type": "FormalParameter",
"dct:conformsTo": "https://bioschemas.org/profiles/FormalParameter/1.0-RELEASE",
"name": "ngpu_config",
"encodingFormat": "text/json",
"valueRequired": true,
"description": "Configuration file of workflow."
"outputs": [
"@type": "FormalParameter",
"dct:conformsTo": "https://bioschemas.org/profiles/FormalParameter/1.0-RELEASE",
"name": "structured_metadata",
"encodingFormat": "text/json",
"description": "Output file of metadata post-processing."
Listing 1: Partial output of an Archivist post-processing pipeline. Extracted information from workflow configuration file was formatted according to the Bioschemas vocabulary.
Details on the hydrology use case
@@
The “Hydrology use case” presents how the metadata Archivist can be applied in the hydrological modeling sciences.
mHM.
@@
An essential tool of this use case is the employed mesoscale hydrologic model mHM [47, 48, 20] (https://mhm-ufz.org, last access: 21 March 2025). mHM is a grid-based, spatially distributed hydrological model driven by daily precipitation, temperature, and potential evapotranspiration. It accounts for major hydrological processes such as snow accumulation and melt, canopy interception, soil infiltration, evapotranspiration, deep percolation, baseflow generation, and river routing. The open-source model code repository is available and is under active development and maintenance (https://git.ufz.de/mhm/mhm, last access: 21 March 2025). A general overview on the model processes and parameterization can be obtained from [47] and [48]. The model is an integral part of the German Drought Monitor (https://www.ufz.de/duerremonitor, last access: 21 March 2025). mHM was also applied and evaluated in multiple climatological regions, including Europe [57, 58], West Africa[59], India[60], and the conterminous United States[61, 50].
Hardware platforms.
@@
mHM simulations were carried out on the Computing Cluster EVE, a joint effort of both the Helmholtz Centre for Environmental Research - UFZ (http://www.ufz.de/, last access: 21 March 2025) and the German Centre for Integrative Biodiversity Research (iDiv) Halle-Jena-Leipzig (https://www.idiv.de/, last access: 25 June 2024). The main compute hardware of EVE comprises a) 42 compute nodes with dual socket Intel Xeon 6348 CPUs with 512 Gigabytes of DDR4 main memory, two of these also include NVIDIA Tesla A100 GPGPUs, and b) 27 compute nodes with dual socket Intel Xeon Gold 6148 CPUs with up to 1536 Gigabytes of DDR4 main memory, two of which include NVIDIA Tesla V100 GPGPUs. The central network component of the cluster is an Intel Omni-Path 100 Series high performance interconnect, providing all compute nodes with non-blocking EDR bandwidth (100 Gigabit per second). All compute nodes share a 4.5 Petabyte IBM Spectrum Scale file system. The system performed with over 164 teraFLOPS under a High-Performance LINPACK (HPL) benchmark. For the simulations of this project, we have used CPU cores exclusively.
Model setup.
@@
The mHM model is executed within the test basin provided along with the model source code, the so called “test basin". The test basin coming with the mHM source code is for the Moselle River basin upstream of Perl, a place, where the Moselle River leaves France and enters Luxembourg and Germany (Moselle Basin). The catchment area is approximately 11,500 $11,500\text{ km}^{2}$, altitude ranges between 150 and 1300 m. a.m.s.l. The Moselle River originates from the Vosges Mountains and is a tributary of the Rhine River. The origin of data used in the test example is provided on https://mhm-ufz.org/docs/ (last access: July 18th, 2024).
Performance metrics.
@@
The hydrology use cases makes use of performance metrics that are commonly used in hydrologic modeling. Explicitly, it is the Kling-Gupta efficiency (KGE [49]), that can be used to compare two time series. The metric combines the long-term mean, long-term variance, and Pearson correlation coefficient.
Data generation.
@@
Data Availability
No external dataset or input data were used for this study.
Code Availability
Source code, example implementations, and additional details of the Archivist framework can be found at https://doi.org/10.5281/zenodo.13442425 (last access: 30 August 2024). An implementation of the metadata post-processing performed for the “Minimal example” can be found at https://doi.org/10.5281/zenodo.13442425 (last access: 30 August 2024). The workflow implementation for the “Neuroscience use case” as well as the data and script required to create fig. 4 can be found at https://doi.org/10.5281/zenodo.13585723 (last access: 30 August 2024). The model implementation of the “Hydrology use case” used to generate the data required to create fig. 5 is available at https://doi.org/10.5281/zenodo.1069202 (last access: 26 January 2025).
Acknowledgments